OpenStudy (anonymous):

Find the derivative using limits of (x^(2)-1)/(2x-3) Using the formula f(x)-f(a)/x-a Thanks!!!

OpenStudy (anonymous):

gotta do it by hand huh?

OpenStudy (anonymous):

this is going to be a raft of algebra we can to it ready?

OpenStudy (anonymous):

lets forget about the \(x-a\) in the denominator, and take care of that last

OpenStudy (anonymous):

\[\frac{x^2-1}{2x-3}-\frac{a^2-1}{2a-3}\] is a start now we actually have to subtract \[\frac{(x^2-1)(2a-3)-(a^2-1)(2x-3)}{(2x-3)(2a-3)}\]

OpenStudy (anonymous):

bunch of algebra at this step, which is why this is really annoying, but when you are done you should end up with \[-\frac{(a-x) (a (2 x-3)-3 x+2))}{(2 a-3) (2 x-3)}\]

OpenStudy (anonymous):

finally divide by \(x-a\) i.e. cancel

OpenStudy (anonymous):

I just got stuck on where to go from (2a-3)(x^(2)-1)-(a^(2)-1)(2x-3) ??

OpenStudy (phi):

the numerator is \[ (2a-3)(x^2-1)-(a^2-1)(2x-3) \] multiply it out to get \[ 2ax^2 -3x^2 -2a + 3 - (2a^2x -3a^2 -2x +3) \\ 2ax^2-3x^2 -2a + 3 -2a^2x + 3a^2 +2x -3 \] now re-arrange and group terms (we are hoping to make form (x-a) factors) \[ (2ax^2-2a^2x) -3x^2+ 3a^2+ (2x-2a) + (3 -3) \\ (2ax^2-2a^2x) - 3(x^2- a^2) + (2x-2a) \] now factor each term \[ 2ax(x-a) - 3(x+a)(x-a) + 2(x-a) \] and finally \[ (2ax -3(x+a) +2) (x-a) \] can you finish?

OpenStudy (anonymous):

Yes!! Thank you. :)